Question 9 Coronado Company has a weighted-average unit contribution margin of $20 for its two products: Drew and Carey. Expected sales for Coronado are 44000 Drews and 56000 Careys. Fixed expenses are $1760000. How many Drews would Coronado sell at the break-even point? 38720 49280 44000 88000
Answer #1

Breakeven point=fixed cost/contribution margin

=(1760000/20)=88000 units

Ratio of sales for Drews:Careys=44000:56000[Total=(44000+56000)=100,000]

Hence target sales for Drews=(44000/100,000)*88000

=38720.